3x/x-3=7+12x^2/x-3
Answer by lwsshak3(11628)

3x/x-3=7+12x^2/x-3
LCD: (x-3)
3x=7x-21+12x^2
12x^2+4x-21=0
(2x+3)(6x-7)=0
2x+3=0
x=-3/2
or
6x-7=0
x=7/6